390 voters are classified by income and political party. The results are shown in the table.

Democrat Republican Total
Low Income 101 75 176
Medium Income 89 63 152
High Income 25 19 44
Super High Income 10 8 18
Total 224 166 390

a. What is the probability of randomly selecting a Democrat in this sample?

b. If a low income voter is randomly chosen, what is the probability that they are a Democrat?

c. Are the events "Democrat" and "Low Income" independent?
Answer #1

a)

probability of randomly selecting a Democrat in this sample =224/390=0.5744

b)

P(democrat given low income voter )=101/176=0.5739

c)

here as P(Democrat) is approximately equal to P(democrat given low income voter ) ; we find them statistically independent.